# Catalog of outbursts of neutron star LMXBs

### v1

@@ Appendix B Frequently outbursting sources
4U 1608-52 2014/01 4.1e-9 5.0e36 Zhang & Yu (2014)
4U 1608-52 2014/10 6.4e-9 7.8e36 Negoro et al. (2014)
4U 1608-52 2016/08 5.5e-9 6.7e36 Ono et al. (2016)
4U 1608-52 2017/06 1.1e-8 1.3e37 Jaisawal et al. (2019)
4U 1608-52 2018/06 8.5e-9 1.0e37 van den Eijnden et al. (2020)
4U 1608-52 2019/03 5.7e-9 7.0e36 MAXI
4U 1608-52 2020/05 1.5e-8 1.8e37 Güver et al. (2021)
4U 1608-52 2021/08 2.8e-9 3.4e36 MAXI
4U 1608-52 2022/07 $>$5.7e-9 $>$7.0e36 MAXI
Table 3: Outbursts of 4U 1608-52. The distance is assumed to be 3.2 kpc. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Spectral fits include $N_{H}=2\times 10^{22}$ cm-2, $\Gamma$=1.7 (Mitsuda et al., 1984) and $N_{H}=1.5\times 10^{22}$, $\Gamma$=1.9 (Gottwald et al., 1987).
MXB 1730-335, aka Rapid Burster, 4U 1730-335, or MXB 1730-33; We use a distance of 8.0$\pm$0.3 kpc, as measured to the host globular cluster, Liller 1 (Baumgardt & Vasiliev, 2021). We include a table for its frequent outbursts (Table B). We note that some of these outbursts may be produced by the second known transient in this cluster, CXOU J173324.6-332321, which was discovered in low-level outburst ($L_{X}$[0.5-10]$=5.5\times 10^{35}$ erg/s) in April 2018 between two larger outbursts (Homan et al., 2018), but does not show Type II bursts as reported from many outbursts (e.g. Bagnoli et al., 2015). We quote MJD values for the highest observed countrate of each outburst. Where only intervals are given (e.g. in 1971), we estimate the midpoint of the interval. For spectral fitting, we assume a 1.6 keV blackbody, with $N_{H}=1.8\times 10^{22}$ cm-2 (Marshall et al., 2001; Mahasena et al., 2003). We use RXTE/PCA Bulge Scan values where available, then RXTE/ASM, MAXI, or Swift/XRT. For RXTE/ASM we multiply countrates by 3.4×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by 2.3×$\times 10^{-12}$. For MAXI we subtract 0.07 cts/cm-2/s (to account for confusion) from the 2-20 keV fluxes, and multiply by 1.1×$\times 10^{-8}$ to get 2-10 keV fluxes. For Swift/XRT data we multiply by 7.5×$\times 10^{-11}$ to get 2-10 keV fluxes, and for INTEGRAL/JEM-X 3-10 keV data we multiply by $1.7\times 10^{-10}$. These different flux estimates are generally in reasonable agreement where they overlap, though the RXTE/ASM has some untrustworthy points when the Sun is very close. We identify 52 outbursts that we do not see specifically identified elsewhere in the literature.
@@
MXB 1730-335 1983/8 45564 1.2e-9 9e36 Kunieda et al. (1984); Barr et al. (1987)
MXB 1730-335 1984/7 45886 1.2e-9 9e36 Kawai et al. (1990)
MXB 1730-335 1985/8-9 46305 2.6e-9 2e37 Stella et al. (1988)
@@
MXB 1730-335 1996/4 50189 6.3e-9 4.8e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1996/10 50387 7.7e-9 5.8e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1997/6 50625 9.3e-9 7.0e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1998/1 50843 6.8e-9 5.2e37 Guerriero et al. (1999); Masetti et al. (2000),RXTE/ASM
MXB 1730-335 1998/8 51045 1.8e-8 1.4e38 Masetti (2002); Mahasena et al. (2003),RXTE/ASM
@@
MXB 1730-335 1999/10 51452 6.8e-9 5.2e37 Masetti (2002); Marshall et al. (2001),RXTE/PCA, RXTE/ASM
MXB 1730-335 2000/3 51633 2.2e-9 1.6e37 Masetti (2002),RXTE/ASM, RXTE/PCA
@@
MXB 1730-335 2011/9 55811 3.6e-9 2.8e37 RXTE/PCA, MAXI, Swift/XRT, IGR/JEM-X
@@
Outbursts of MXB 1730-335, the Rapid Burster in the globular cluster Liller 1. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Distance of 8 kpc assumed.
Note. —
GRS 1747-312; We use the distance estimate for the globular cluster Terzan 6 of 7.3±$\pm 0.4$ kpc from Baumgardt & Vasiliev (2021). GRS 1747-312’s outbursts have not been catalogd since 2008, so we turn to all-sky monitors to produce a full list (Table B). We assume $N_{H}=1.4\times 10^{22}$ cm-2 and a photon index of 1.0 (in’t Zand et al., 2000) to transform countrates to fluxes. For RXTE/ASM we multiply countrates by 3.6×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by $\times 10^{-12}$. For Swift/XRT data we multiply by 9.4×$\times 10^{-11}$ to get 2-10 keV fluxes, for MAXI by $8.7\times 10^{-9}$, and for INTEGRAL/JEM-X by $7.1\times 10^{-11}$. We identify outbursts in the RXTE/ASM, RXTE/PCA, INTEGRAL/JEM-X, and MAXI lightcurves when we see at least two detections (among all the lightcurves) with the countrate exceeding 3 times the error (see Fig. 9). This generally agrees with the identifications in in’t Zand et al. (2003) where they overlap, though the ASM data is noisy in this crowded region. We are also in general agreement with the identifications of Šimon (2009), but are not completely confident of the reality of their epochs 13, 20, and 21 (not included in our list), as the signal-to-noise of the individual detections is not above 3 sigma per day, though the timing of the outbursts appears plausible. The outbursts average every 138 days, with the average varying between 120 and 150 days over time, consistent with previous determinations (in’t Zand et al., 2003; Šimon, 2009). Peak fluxes average $L_{X}=6\times 10^{36}$ erg/s.
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.111111van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X-2” in Table B, and Fig. 9.
@@
GRS 1747-312 1996/3 0 50158 1.6e-9 1.0e37 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1996/8 1 50319 1.3e-9 8.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/1 2 50474 1.2e-9 7.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/5 3 50598 1.3e-9 8.5e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/10 4 50713 1.3e-9 8.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/2 5 50865 1.1e-9 7.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/6 6 50985 6.4e-10 4.0e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312* 1998/9 7 51058 1.0e-9 6.4e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1999/1 8 51189 3.0e-9 1.9e37 Šimon (2009), RXTE/ASM
GRS 1747-312 1999/5 9 51308 1.1e-9 5.9e36 in’t Zand et al. (2000, 2003); Šimon (2009), PCA, RXTE/ASM
GRS 1747-312* 1999/10 10 51448 8.3e-10 5.3e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/2 11 51586 8.0e-10 5.1e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/8 12 51760 1.4e-9 8.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/5 14 52062 1.2e-9 7.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/10 15 52209 1.0e-9 6.5e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/2 16 52331 8.1e-10 5.2e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/7 17 52468 8.9e-10 5.7e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2002/10 18 52574 8.8e-10 5.6e36 Šimon (2009), RXTE/ASM
GRS 1747-312 2003/3 19 52713 1.2e-9 7.6e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/3 22 53083 7.2e-10 4.5e36 Markwardt & Swank (2004a); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/8 23 53233 8.5e-10 5.4e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/11 24 53341 9.7e-10 6.1e36 Šimon (2009), RXTE/ASM
GRS 1747-312 2005/4 25 53482 7.8e-10 4.9e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2005/9 26 53624 5.8e-10 3.7e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/2 27 53775 8.5e-10 5.4e36 Chenevez et al. (2006); Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/7 28 53923 9.5e-10 6.0e36 Šimon (2009) RXTE/ASM, PCA
GRS 1747-312 2007/1 29 54129 1.3e-9 8.2e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/6 30 54265 1.4e-9 8.7e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/11 31 54404 8.0e-10 5.1e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2008/3 32 54556 6.9e-10 4.4e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2008/9 54726 8.3e-10 5.3e36 RXTE/ASM, IGR/JEM-X
GRS 1747-312 2009/2 54869 1.2e-9 7.8e36 RXTE/ASM, PCA
GRS 1747-312 2009/6 55008 1.2e-9 7.7e36 RXTE/ASM, PCA
Ter 6 X-2 2009/9 55090 1.0e-11 6e34 Saji et al. (2016); van den Berg et al. (2024)
GRS 1747-312 2009/11 55145 1.2e-9 7.7e36 RXTE/ASM, MAXI
@@
GRS 1747-312* 2013/3 56369 7.9e-10 5.0e36 Chenevez et al. (2013); Bahramian et al. (2013a), Swift/XRT, MAXI, JEM-X

### v2

@@ Appendix B Frequently outbursting sources
4U 1608-52 2014/1 4.1e-9 5.0e36 Zhang & Yu (2014)
4U 1608-52 2014/10 6.4e-9 7.8e36 Negoro et al. (2014b)
4U 1608-52 2016/8 5.5e-9 6.7e36 Ono et al. (2016)
4U 1608-52 2017/6 1.1e-8 1.3e37 Jaisawal et al. (2019)
4U 1608-52 2018/6 8.5e-9 1.0e37 van den Eijnden et al. (2020)
4U 1608-52 2019/3 5.7e-9 7.0e36 MAXI
4U 1608-52 2020/5 1.5e-8 1.8e37 Güver et al. (2021)
4U 1608-52 2021/8 2.8e-9 3.4e36 MAXI
4U 1608-52 2022/7-2023/7 2.1e-8 2.6e37 Chen et al. (2023b, 2024), MAXI
4U 1608-52 2024/2 4e-9 5e36 MAXI
4U 1608-52 2025/3 1.8e-8 2.2e37 Wang et al. (2025),MAXI
Table 3: Outbursts of 4U 1608-52. The distance is assumed to be 3.2 kpc. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Spectral fits assume $N_{H}=2\times 10^{22}$ cm-2, $\Gamma$=1.7 (Mitsuda et al., 1984) and $N_{H}=1.5\times 10^{22}$, $\Gamma$=1.9 (Gottwald et al., 1987). Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
MXB 1730-335, aka Rapid Burster, 4U 1730-335, or MXB 1730-33. We use a distance of 8.0$\pm$0.3 kpc, as measured to the host globular cluster, Liller 1 (Baumgardt & Vasiliev, 2021). We include a table for its frequent outbursts (Table B). We note that some of these outbursts may be produced by the second known transient in this cluster, CXOU J173324.6-332321, which was discovered in low-level outburst ($L_{X}$[0.5-10]$=5.5\times 10^{35}$ erg/s) in April 2018 between two larger outbursts (Homan et al., 2018), but likely does not show Type II bursts as reported from many outbursts (e.g. Bagnoli et al., 2015). We quote MJD values for the highest observed countrate of each outburst. Where only intervals are given (e.g. in 1971), we estimate the midpoint of the interval. For spectral fitting, we assume a 1.6 keV blackbody, with $N_{H}=1.8\times 10^{22}$ cm-2 (Marshall et al., 2001; Mahasena et al., 2003). We use RXTE/PCA Bulge Scan values where available, then RXTE/ASM, MAXI, or Swift/XRT. For RXTE/ASM we multiply countrates by 3.4×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by 2.3×$\times 10^{-12}$. For MAXI we subtract 0.07 cts/cm-2/s (to account for confusion) from the 2-20 keV fluxes, and multiply by 1.1×$\times 10^{-8}$ to get 2-10 keV fluxes. For Swift/XRT data we multiply by 7.5×$\times 10^{-11}$ to get 2-10 keV fluxes, and for INTEGRAL/JEM-X 3-10 keV data we multiply by $1.7\times 10^{-10}$. These different flux estimates are generally in reasonable agreement where they overlap, though the RXTE/ASM data has some untrustworthy points when the Sun is very close. We identify ******* 52 outbursts that we do not see specifically identified elsewhere in the literature.
@@
MXB 1730-335 1983/8 45564 1.2e-9 9e36 Tanaka (1983); Kunieda et al. (1984); Barr et al. (1987)
MXB 1730-335 1984/7 45886 1.2e-9 9e36 Makino (1984); Kawai et al. (1990)
MXB 1730-335 1985/8-9 46305 2.6e-9 2e37 Stella et al. (1985, 1988)
@@
MXB 1730-335 1996/4 50189 6.3e-9 4.8e37 Levine et al. (1996b); Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1996/10 50387 7.7e-9 5.8e37 Remillard (1996); Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1997/6 50625 9.3e-9 7.0e37 Guerriero et al. (1997, 1999),RXTE/ASM
MXB 1730-335 1998/1 50843 6.8e-9 5.2e37 Guerriero et al. (1998, 1999); Masetti et al. (2000),RXTE/ASM
MXB 1730-335 1998/8 51045 1.8e-8 1.4e38 Fox et al. (1998); Masetti (2002); Mahasena et al. (2003),RXTE/ASM
@@
MXB 1730-335 1999/10 51452 6.8e-9 5.2e37 Fox & Lewin (1999); Masetti (2002); Marshall et al. (2001),RXTE/PCA,ASM
MXB 1730-335 2000/3 51633 2.2e-9 1.6e37 Fox & Lewin (2000); Masetti (2002),RXTE/ASM, RXTE/PCA
@@
MXB 1730-335 2011/9 55811 3.6e-9 2.8e37 Chenevez et al. (2011a), RXTE/PCA, MAXI, Swift/XRT, IGR/JEM-X
@@
MXB 1730-335 2024/1 60329 1.9e-9 1.4e37 MAXI
MXB 1730-335 2024/6 60490 2.6e-9 2.0e37 MAXI
MXB 1730-335 2024/10 60594 2.1e-9 1.6e37 MAXI
MXB 1730-335 2025/1 60689 1.3e-9 9.5e36 MAXI
MXB 1730-335 2025/4 60791 1.0e-9 7.5e36 MAXI
Note. — Outbursts of MXB 1730-335, the Rapid Burster in the globular cluster Liller 1. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Distance of 8 kpc assumed. Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
GRS 1747-312; We use the distance estimate for the globular cluster Terzan 6 of 7.3±$\pm 0.4$ kpc from Baumgardt & Vasiliev (2021). GRS 1747-312’s outbursts have not been cataloged since 2008, so we turn to all-sky monitors to produce a full list (Table B). We assume $N_{H}=1.4\times 10^{22}$ cm-2 and a photon index of 1.0 (in’t Zand et al., 2000) to transform countrates to fluxes. For RXTE/ASM we multiply countrates by 3.6×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by $\times 10^{-12}$. For Swift/XRT data we multiply by 9.4×$\times 10^{-11}$ to get 2-10 keV fluxes, for MAXI by $8.7\times 10^{-9}$, and for INTEGRAL/JEM-X by $7.1\times 10^{-11}$. We average RXTE/ASM data over a 3-day window (weighting by the error), to reduce the noise. We identify outbursts in the RXTE/ASM, RXTE/PCA, INTEGRAL/JEM-X, and MAXI lightcurves when we see at least two detections (among all the lightcurves) with the countrate exceeding 3 times the error above the average quiescent rate of 0.5 ct/s (see Fig. 10). This generally agrees with the identifications in in’t Zand et al. (2003) where they overlap, though the ASM data is noisy in this crowded region. We are also in good agreement with the identifications of Šimon (2009), which use a more complicated smoothing algorithm to construct their lightcurves. We note that the last two outbursts detected by the RXTE/ASM (late 2010 to early 2011) are especially noisy; we do not use these data, though we plot them in Fig. 10. The outbursts average every 138 days, with the average varying between 120 and 150 days over time, consistent with previous determinations (in’t Zand et al., 2003; Šimon, 2009). Peak fluxes average $L_{X}=6\times 10^{36}$ erg/s.
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.121212van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X2” in Table B.
@@
GRS 1747-312 1996/3 0 50157 1.3e-9 8.2e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1996/8 1 50318 1.1e-9 6.8e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/1 2 50475 9.4e-10 6.0e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/5 3 50598 1.2e-9 7.4e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/10 4 50714 1.1e-9 7.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/2 5 50867 1.2e-9 7.5e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/6 6 50986 3.9e-10 2.5e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312* 1998/9 7 51057 9.1e-10 5.8e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1999/1 (8) 51189 1.7e-9 1.1e37 Šimon (2009), RXTE/ASM
GRS 1747-312 1999/5 9 51308 1.0e-9 6.3e36 in’t Zand et al. (2000, 2003); Šimon (2009), PCA, RXTE/ASM
GRS 1747-312* 1999/10 10 51451 9.9e-10 6.3e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/2 11 51585 1.3e-9 8.2e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/8 12 51759 1.1e-9 6.9e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/5 14 52063 9.5e-10 6.0e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/10 15 52207 1.0e-9 6.7e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/2 16 52337 8.8e-10 5.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/7 17 52470 9.6e-10 6.1e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2003/3 19 52713 8.7e-10 5.5e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/3 22 53100 1.0e-9 6.5e36 Markwardt & Swank (2004a); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/8 23 53233 1.1e-9 6.7e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/11 24 53341 1.0e-9 6.4e36 Šimon (2009), RXTE/ASM
GRS 1747-312 2005/4 25 53489 9.1e-10 5.8e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2005/9 26 53633 8.8e-10 5.6e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/2 27 53782 1.1e-9 6.9e36 Chenevez et al. (2006); Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/7 28 53926 1.4e-9 8.8e36 Šimon (2009) RXTE/ASM, PCA
GRS 1747-312 2007/1 29 54128 1.4e-9 8.9e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/6 30 54264 1.2e-9 7.9e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/11 31 54406 9.1e-10 5.8e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2008/3 32 54552 7.5e-10 4.8e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2008/9 54726 5.9e-10 3.7e36 IGR/JEM-X
GRS 1747-312 2009/2 54870 1.1e-9 7.1e36 RXTE/ASM, PCA
GRS 1747-312 2009/6 55008 1.1e-9 7.0e36 RXTE/ASM, PCA
Ter 6 X2 2009/9 55090 1.0e-11 6e34 Saji et al. (2016); van den Berg et al. (2024)
GRS 1747-312 2009/11 55147 1.1e-9 7.0e36 RXTE/ASM, MAXI
@@
GRS 1747-312* 2013/3 56369 7.9e-10 5.0e36 Chenevez et al. (2013b); Bahramian et al. (2013a), Swift/XRT, MAXI, JEM-X